# VideoAgent: Self-Improving Video Generation

### v2

@@ Appendix F VLM Feedback for Correction
VideoAgent 46.7% 29.3% 13.3% 9.3% 38.7% 18.7%
@@
Appendix G Details of Human Evaluation on BridgeData V2
@@
Next, we qualitatively evaluate video generation quality using the five Video-Score dimensions: Visual Quality (VQ) for clarity and resolution, Temporal Consistency (TC) for smooth frame transitions, Dynamic Degree (DD) for capturing accurate object/environment changes, Text-to-Video Alignment (TVA) for matching the video to the prompt, and Factual Consistency (FC) for adherence to physical laws and real-world facts. Videos are rated on a 4-point scale based on the metric in He et al. (2024): 1 (Bad), 2 (Average), 3 (Good), and 4 (Perfect). Our evaluation is based on 50 generated videos from a held-out set.
@@
Appendix H Examples
H.1 Improvements in Meta-World
@@ H.2 Improvements in iThor
H.3 Identification and Suggestive Feedback Examples

### v3

@@ E.3 Further Analysis of VideoAgent-Online
VideoAgent 40.0% 29.3% 13.3% 9.3% 38.7% 18.7%
VideoAgent-Online (Iter1) 48.0% 40.0% 24.0% 12.0% 42.7% 36.0%
VideoAgent-Online (Iter2) 58.7% 50.7% 28.0% 18.7% 53.3% 41.3%
VideoAgent-Online (Iter3) 58.7% 52.0% 26.7% 17.3% 54.7% 40.0%
faucet-close faucet-open handle-press hammer assembly Overall
AVDC 12.0% 17.3% 41.3% 0.00% 5.30% 19.6%
VideoAgent 46.7% 12.0% 36.0% 0.0% 1.3% 22.3%
VideoAgent-Online (Iter1) 53.3% 28.0% 52.0% 1.3% 5.3% 31.2%
VideoAgent-Online (Iter2) 58.7% 36.0% 64.0% 1.3% 9.3% 38.2%
VideoAgent-Online (Iter3) 56.3% 36.0% 66.7% 1.3% 10.7% 38.22%
Table 7: Meta-World Result. The mean success rates of VideoAgent combined with successive rounds of data collection via Online Iterations and Replan modules as compared to AVDC baseline.
Appendix F Architectural Details of VideoAgent
F.1 Video Diffusion training details
F.2 Inference time speed
In our current setup, during inference, our video generation model produces a new video within 10 seconds on a single A6000 GPU at a resolution of $128\times 128$ for Meta-World. The process of mapping this generated video to an action takes, on average, an additional 25 seconds. This action-mapping stage involves calculating optical flow, receiving feedback from the vision-language model (VLM), and converting the video into an action sequence based on the computed flow.
Appendix G VLM Feedback for Correction
AVDC 30.7% 28.0% 21.3% 8.00% 34.7% 17.3%
VideoAgent 40.0% 29.3% 13.3% 9.3% 38.7% 18.7%
@@ Appendix G VLM Feedback for Correction
Appendix H Details of Human Evaluation on BridgeData V2
@@
Next, we qualitatively evaluate video generation quality using the five Video-Score dimensions: Visual Quality (VQ) for clarity and resolution, Temporal Consistency (TC) for smooth frame transitions, Dynamic Degree (DD) for capturing accurate object/environment changes, Text-to-Video Alignment (TVA) for matching the video to the prompt, and Factual Consistency (FC) for adherence to physical laws and real-world facts. Videos are rated on a 4-point scale based on the metric in (He et al., 2024): 1 (Bad), 2 (Average), 3 (Good), and 4 (Perfect). Our evaluation is based on 50 generated videos from a held-out set.
@@
Appendix I Examples
I.1 Zero-shot generalization on real-world scenes
VideoAgent trained on Bridge dataset demonstrates strong performance on zero shot video generation for natural distribution shifts and longer language instructions. Some examples of the synthesized videos can be found in Fig. 8.
Figure 8: Zero-shot generalization of VideoAgent: VideoAgent generalizes fairly well to natural distribution shifts and is able to generate successful trajectories on data it has not been trained on.
I.2 Improvements in Meta-World
@@ I.3 Improvements in iThor
I.4 Identification and Suggestive Feedback Examples